# A lower bound for the Weisfeiler-Leman dimension of circulant graphs

### v1

@@ 10.1. Reduction from S-rings to graphs
Let ${\cal G}\in{\mathfrak{B}}_{a}$ be an $\varepsilon$-expander for some real $\varepsilon>0$. This means that for every vertex subset $\Delta$ containing at most half of the vertices of ${\cal G}$, we have $|\partial\Delta|\geq\varepsilon|\Delta|$, where $\partial\Delta$ is the set of all vertices outside $\Delta$, each adjacent to at least one vertex of $\Delta$. It is not hard to prove (see, e.g., [5, Lemma 4.2.7]) that each separator of the graph ${\cal G}$ has cardinality greater than $k=\lfloor c_{\varepsilon}\cdot 2a\rfloor+1$, where $c_{\varepsilon}={\varepsilon\over{4+\varepsilon}}$.
Let ${\cal A}^{\star}$ be the totally normal S-ring over a cyclic group $G$ of order $n$ and $\varphi^{\star}=\varphi_{\mathfrak{c}}$ the algebraic isomorphism of ${\cal A}^{\star}$, that were defined in Theorem 9.1; in particular, ${\cal A}^{\star}$ admits a $\lfloor k/6\rfloor$-local system of inner ${\cal A}^{\star}$-multipliers with respect to ${\mathfrak{c}}$ and $\operatorname{Iso}({\cal A}^{\star},\varphi^{\star})=\varnothing$. Then
(53) $\operatorname{Iso}({\cal X}^{\star},\varphi^{\star})=\varnothing,$
where ${\cal X}^{\star}={\cal X}({\cal A}^{\star})$. Furthermore, from Theorem 7.1, it follows that ${\cal X}^{\star}$ is $\operatorname{WL}_{m}$-equivalent to itself with respect to the algebraic isomorphism $\varphi^{\star}$, where $m=\lfloor\sqrt{\lfloor k/6\rfloor}-1\rfloor=c\cdot\sqrt{a}$ for a certain real constant $10<c<$ 1. This fact together with relation (53) shows that
(54) $\dim_{\scriptscriptstyle\operatorname{WL}}({\cal X}^{\star})\geq c\cdot\sqrt{a},$
see Section 4. In the rest of this section, we construct a circulant graph ${\cal G}^{\star}$ over the group $G$ and verify that
(55) $\operatorname{WL}({\cal G}^{\star})={\cal X}^{\star}.$
This is enough to complete the proof. Indeed, in accordance with formulas (9) and (54), we then have $\dim_{\scriptscriptstyle\operatorname{WL}}({\cal G}^{\star})=\dim_{\scriptscriptstyle\operatorname{WL}}({\cal X}^{\star})\geq c\cdot\sqrt{a}$, as required.
10.2. The definition of the graph ${\cal G}^{\star}$
To construct the graph ${\cal G}^{\star}$, for any two indices $1\leq i,j\leq b$, we choose an arbitrary basic set $X_{i,j}\in{\cal S}({\cal A}^{\star})$ contained in the difference $D_{i,j}$ defined by formula (45) and such that $(X_{i,j})_{{\mathfrak{s}}_{i,j}}$ is a highest basic set of the S-ring ${\cal A}_{{\mathfrak{s}}_{i,j}}$, where ${\mathfrak{s}}_{i,j}:={\mathfrak{s}}_{u_{i},w_{j}}$. We have
(56) $|X_{i,j}|=|\pi_{i,j}^{-1}((X_{i,j})_{{\mathfrak{s}}_{i,j}})|=|G_{i-1,j-1}|\cdot\begin{cases}4&\text{if $i=1$ or $j=1$},\\ 8&\text{if $i,j>1$ and $\{u_{i},w_{j}\}\in E$},\\ 16&\text{otherwise},\\ \end{cases}$
@@ 10.2. The definition of the graph ${\cal G}^{\star}$
${\cal G}^{\star}=\operatorname{Cay}(G,X^{\star})$
@@ 10.3. Proof of equality (55)
The scheme $\operatorname{WL}({\cal G}^{\star})$ is a Cayley scheme over the group $G$. It is associated with the S-ring ${\cal A}^{\prime}=\operatorname{WL}_{G}({\underline{X^{\star}}})$. Since $X^{\star}$ is also an ${\cal A}^{\star}$-set, we have ${\cal A}^{\star}\geq{\cal A}^{\prime}$. It remains to prove the reverse inclusion
(57) ${\cal A}^{\star}\leq{\cal A}^{\prime},$
i.e., that each basic set of ${\cal A}^{\star}$ is an ${\cal A}^{\prime}$-set. To do this, we need several auxiliary statements.
@@
We prove the statement for $G_{u_{b}}$ only; the statement for $G_{w_{b}}$ is proved similarly. Recall that the number $n_{1}$ is a product of two distinct primes, say $p_{1}$ and $q_{1}$. We apply Theorem 5.1(2) for ${\cal A}={\cal A}^{\prime}$ and $X=X^{\star}$. Note that if $x\in X^{\star}$ and $H$ the subgroup defined before that theorem, then $xH\subseteq X^{\star}$ unless $x$ belongs to the union of all $X_{i,j}$ with $i=0,1$. Therefore,
@@
where for all $i,j$, we set $X^{\prime}_{i,j}=\sum_{x\in X_{i,j}}x^{p_{1}}$ and $X^{\prime\prime}_{i,j}=\sum_{x\in X^{\prime}_{i,j}}x^{q_{1}}$. The right-hand side is contained in $G_{u_{b}}$ and contains the set $X^{\prime\prime}_{0,b}$ generating $G_{u_{b}}$. Thus, $G_{u_{b}}$ is an ${\cal A}^{\prime}$-group. ∎
@@
From Claim 2, it follows that $X:=X^{\star}\cap G_{u_{b}}$ is an ${\cal A}^{\prime}$-set. Note that $X=X_{0,0}\cup X_{1,0}\cup\ldots\cup X_{b,0}$ and $X_{i,0}\subseteq D_{i,0}$ for $i=0,1,\ldots,b$. Since each $X_{i,0}$ is a basic set of the S-ring ${\cal A}^{\star}$, we have
@@
Besides, $a_{0}=|X|$. Together with formula (56), this shows that the integers $a_{0},a_{1}.\ldots,a_{b}$ are pairwise distinct. It easily follows that $X_{i,0}$ is an ${\cal A}^{\prime}$-set for all $i$. Therefore, $G_{u_{i}}=\langle{X_{i,0}}\rangle$ is an ${\cal A}^{\prime}$-group, $i=1,\ldots,b$. In a similar way, one can prove that so is the group $G_{w_{j}}=\langle{X_{0,j}}\rangle$, $j=1,\ldots,b$. ∎
@@
Let $10\leq i,j\leq b-$ 1. Then ${\mathfrak{s}}_{i,j}\in{\mathfrak{S}}({\cal A}^{\prime})$ and $D_{i,j},X_{i,j}\in{\cal S}({\cal A}^{\prime})^{\cup}$.
@@
By Claim 3, $G_{i,j}=G_{u_{i}}\times G_{w_{j}}$ is an ${\cal A}^{\prime}$-group for all $i,j$. Then obviously ${\mathfrak{s}}_{i,j}$ is an ${\cal A}^{\prime}$-section and $D_{i,j}$ is an ${\cal A}^{\prime}$-set. Moreover, so is the subset $X_{i,j}=X^{\star}\cap D_{i,j}$. ∎
Let $X$ be a nonidentity basic set of ${\cal A}^{\star}$. Then there are indices $10\leq i,j\leq b-$ 1 such that $X\subseteq D_{i,j}$. Put ${\mathfrak{s}}={\mathfrak{s}}_{i,j}$. From Lemma 9.3, it follows that $({\cal A}^{\star})_{\mathfrak{s}}$ is a normal S-ring over ${\mathfrak{s}}$. By the choice of the basic set $X_{i,j}$, Lemma 5.2 implies that $({\cal A}^{\star})_{\mathfrak{s}}$ is the smallest S-ring over ${\mathfrak{s}}$ that contains $(X_{i,j})_{\mathfrak{s}}$. By Corollary 10.1, this implies that $({\cal A}^{\prime})_{\mathfrak{s}}\geq({\cal A}^{\star})_{\mathfrak{s}}$. Therefore,
$X_{\mathfrak{s}}\in{\cal S}({\cal A}^{\star}_{\mathfrak{s}})^{\cup}\subseteq{\cal S}({\cal A}^{\prime}_{\mathfrak{s}})^{\cup}.$
By formula (46), this shows that $X=\pi_{i,j}^{-1}(X_{\mathfrak{s}})$ is an ${\cal A}^{\prime}$-set, which proves inclusion (57).
@@
Recall (see, e.g., [20] and references therein) that a $k$-regular graph ${\cal G}$ is Ramanujan if all of its nontrivial eigenvalues lie between $-2\sqrt{k-1}$ and $2\sqrt{k-1}$. Any such graph is an $\varepsilon$-expander with $2\varepsilon=1-\lambda/k$, where $\lambda$ is the second largest (in absolute value) eigenvalue of the adjacency matrix of ${\cal G}$. By Theorem 1.1, it suffices to find an infinite family of bipartite cubic $3$-connected Ramanujan graphs.
Let $d\in{\mathbb{N}}$ be even. Denote by ${\cal G}$ a connected cubic Ramanujan graph on $2^{3d}-2^{d}$ vertices, constructed in [22, Theorem 5.13]. Note that it is a Cayley graph of $\operatorname{PSL}(2,2^{d})$. Let ${\cal G}_{d}$ be the $2$-cover of ${\cal G}$, i.e., the vertex set of ${\cal G}_{d}$ is the disjoint union of the vertex sets of two copies of ${\cal G}$ and two vertices of ${\cal G}_{d}$ are adjacent if they lie in different copies and the corresponding vertices of ${\cal G}$ are adjacent. Obviously, the graph ${\cal G}_{d}$ is cubic and bipartite. Moreover, it is a Ramanujan graph (see, e.g., a remark in [21, p. 311]).
It remains to verify that the graph ${\cal G}_{d}$ is $3$-connected. From the fact that ${\cal G}$ is a Cayley graph, it easily follows that the graph ${\cal G}_{d}$ is vertex transitive. The connectivity of it can be equal neither $1$ (because it contains at least one vertex which is not a cut point), nor $2$ (because any vertex transitive graph of connectivity $2$ is a polygon [29, Corollary 3A]). Thus the (cubic) graph ${\cal G}_{d}$ is $3$-connected.

### v2

@@ 10.1. Reduction from S-rings to graphs
Let ${\mathcal{G}}\in{\mathfrak{B}}_{a}$ be an $\varepsilon$-expander for some real $\varepsilon>0$. This means that for every vertex subset $\Delta$ containing at most half of the vertices of $\mathcal{G}$}, we have $|\partial\Delta|\geq\varepsilon|\Delta|$, where $\partial\Delta$ is the set of all vertices outside $\Delta$, each adjacent to at least one vertex of $\Delta$. It is not hard to prove (see, e.g., [5, Lemma 4.2.7]) that each separator of the graph $\mathcal{G}$} has cardinality greater than $k=\lfloor c_{\varepsilon}\cdot 2a\rfloor+1$, where $c_{\varepsilon}={\varepsilon\over{4+\varepsilon}}$.
[…]
(53) $\operatorname{Iso}({\mathcal{X}}^{\star},\varphi^{\star})=\varnothing,$
where ${\mathcal{X}}^{\star}={\mathcal{X}}({\mathcal{A}}^{\star})$. Furthermore, from Theorem 7.1, it follows that $\mathcal{X}$$^{\star}$ is $\operatorname{WL}_{m}$-equivalent to itself with respect to the algebraic isomorphism $\varphi^{\star}$, where $m=\lfloor\sqrt{\lfloor k/6\rfloor}-1\rfloor=c\cdot\sqrt{a}$ for a certain real constant $0<c<1$. This fact together with relation (53) shows that
(54) $\dim_{\scriptscriptstyle\operatorname{WL}}({\mathcal{X}}^{\star})\geq c\cdot\sqrt{a},$
see Section 4. In the rest of this section, we construct a circulant graph $\mathcal{G}$$^{\star}$ over the group $G$ and verify that
(55) $\operatorname{WL}({\mathcal{G}}^{\star})={\mathcal{X}}^{\star}.$
This is enough to complete the proof. Indeed, in accordance with formulas (9) and (54), we then have $\dim_{\scriptscriptstyle\operatorname{WL}}({\mathcal{G}}^{\star})=\dim_{\scriptscriptstyle\operatorname{WL}}({\mathcal{X}}^{\star})\geq c\cdot\sqrt{a}$, as required.
10.2. The definition of the graph $\mathcal{G}$$^{\star}$
To construct the graph $\mathcal{G}$$^{\star}$, for any two indices $1\leq i,j\leq b$, we choose an arbitrary basic set $X_{i,j}\in{\mathcal{S}}({\mathcal{A}}^{\star})$ contained in the difference $D_{i,j}$ defined by formula (45) and such that $(X_{i,j})_{{\mathfrak{s}}_{i,j}}$ is a highest basic set of the S-ring ${\mathcal{A}}_{{\mathfrak{s}}_{i,j}}$, where ${\mathfrak{s}}_{i,j}:={\mathfrak{s}}_{u_{i},w_{j}}$. We have
[…]
@@ 10.2. The definition of the graph $\mathcal{G}$$^{\star}$
${\mathcal{G}}^{\star}=\operatorname{Cay}(G,X^{\star})$
@@ 10.3. Proof of equality (55)
[…]
(57) ${\mathcal{A}}^{\star}\leq{\mathcal{A}}^{\prime},$
[…]
@@
We prove the statement for $G_{u_{b}}$ only; the statement for $G_{w_{b}}$ is proved similarly. Recall that the number $n_{1}$ is a product of two distinct primes, say $p_{1}$ and $q_{1}$. We apply Theorem 5.1(2) for ${\mathcal{A}}={\mathcal{A}}^{\prime}$ and $X=X^{\star}$. Note that if $x\in X^{\star}$ and $H$ the subgroup defined before that theorem, then $xH\subseteq X^{\star}$ unless $x$ belongs to the union of all $X_{i,j}$ with $i=0,1$. Therefore,
@@
[…]
@@
[…]
@@
[…]
@@
Let $0\leq i,j\leq b-1$. Then ${\mathfrak{s}}_{i,j}\in{\mathfrak{S}}({\mathcal{A}}^{\prime})$ and $D_{i,j},X_{i,j}\in{\mathcal{S}}({\mathcal{A}}^{\prime})^{\cup}$.
@@
[…]
Let $X$ be a nonidentity basic set of $\mathcal{A}$$^{\star}$. Then there are indices $0\leq i,j\leq b-1$ such that $X\subseteq D_{i,j}$. Put ${\mathfrak{s}}={\mathfrak{s}}_{i,j}$. From Lemma 9.3, it follows that $({\mathcal{A}}^{\star})_{\mathfrak{s}}$ is a normal S-ring over $\mathfrak{s}$}. By the choice of the basic set $X_{i,j}$, Lemma 5.2 implies that $({\mathcal{A}}^{\star})_{\mathfrak{s}}$ is the smallest S-ring over $\mathfrak{s}$} that contains $(X_{i,j})_{\mathfrak{s}}$. By Corollary 10.1, this implies that $({\mathcal{A}}^{\prime})_{\mathfrak{s}}\geq({\mathcal{A}}^{\star})_{\mathfrak{s}}$. Therefore,
$X_{\mathfrak{s}}\in{\mathcal{S}}({\mathcal{A}}^{\star}_{\mathfrak{s}})^{\cup}\subseteq{\mathcal{S}}({\mathcal{A}}^{\prime}_{\mathfrak{s}})^{\cup}.$
[…]
@@
Recall (see, e.g., [20] and references therein) that a $k$-regular graph $\mathcal{G}$} is Ramanujan if all of its nontrivial eigenvalues lie between $2\sqrt{k-1}$ and $2\sqrt{k-1}$. Any such graph is an $\varepsilon$-expander with $2\varepsilon=1-\lambda/k$, where $\lambda$ is the second largest (in absolute value) eigenvalue of the adjacency matrix of $\mathcal{G}$}. By Theorem 1.1, it suffices to find an infinite family of bipartite cubic $3$-connected Ramanujan graphs.
Let $d\in{\mathbb{N}}$ be even. Denote by $\mathcal{G}$} a connected cubic Ramanujan graph on $2^{3d}-2^{d}$ vertices, constructed in [22, Theorem 5.13]. Note that it is a Cayley graph of $\operatorname{PSL}(2,2^{d})$. Let ${\mathcal{G}}_{d}$ be the $2$-cover of $\mathcal{G}$}, i.e., the vertex set of ${\mathcal{G}}_{d}$ is the disjoint union of the vertex sets of two copies of $\mathcal{G}$} and two vertices of ${\mathcal{G}}_{d}$ are adjacent if they lie in different copies and the corresponding vertices of $\mathcal{G}$} are adjacent. Obviously, the graph ${\mathcal{G}}_{d}$ is cubic and bipartite. Moreover, it is a Ramanujan graph (see, e.g., a remark in [21, p. 311]).
It remains to verify that the graph ${\mathcal{G}}_{d}$ is $3$-connected. From the fact that $\mathcal{G}$} is a Cayley graph, it easily follows that the graph ${\mathcal{G}}_{d}$ is vertex transitive. The connectivity of it can be equal neither $1$ (because it contains at least one vertex which is not a cut point), nor $2$ (because any vertex transitive graph of connectivity $2$ is a polygon [29, Corollary 3A]). Thus the (cubic) graph ${\mathcal{G}}_{d}$ is $3$-connected.
Acknowledgment
The authors would like to thank Gang Chen for his useful discussion and helpful suggestions.